To prevent dehydration, intravenous (IV) solution is given to the hospital patients. Commonly used are solutions containing sodium chloride and glucose. A 150.0ml IV solution contains 6.60g glucose (C6H1206). What is the molarity of this solution?


1
Expert's answer
2021-03-03T06:29:43-0500

n(glucose)= 6.60/180 = 0.0367 mol

C = n/V = 0.0367/0.15 = 0.244 mol/L

Answer: 0.244 mol/L